OpenStudy (anonymous):

Given the recursively defined sequence, find a_4 a_1 = -3 an = 2n – a_(n-1)

OpenStudy (anonymous):

That should be a_n

OpenStudy (anonymous):

\(a_1=-3\) so \(a_2=2\times 2-(-3)=4+3=7\)

OpenStudy (anonymous):

you got this or no?

OpenStudy (anonymous):

Yeah ty.

OpenStudy (anonymous):

yw so you can get \(a_3\) and \(a_4\) the same way, right?

OpenStudy (anonymous):

Yep